The number of moles of KMnO4 that will be needed to react completely with one mole of ferrous oxalate inn acidic medium is
a. 2/5
b. 3/5
c. 4/5
d. 1

Grade:upto college level

3 Answers

Deepak Patra
9 years ago
Sol. The balanced chemical reaction is : 3MnO_4^- + 5 FeC2O¬2 + 24H+ → 3Mn2+ + 5Fe3+ + 10CO2 + 12H2O ∵ 5 moles of FeC2O4 reacts with 3 moles of KMnO4 ∴ 1 mole of FeC2O4 will require 3/5 mole of KMnO4.
Gaurav
5 years ago
No. Of equivalents of kmno4=no. Of equivalents of fec2o4
Equiv. Of kmno4=n × no. Of moles=5× x.
Equivalents of fec2o4= 3×1=3
5 × x=3
x=3/5 answer.
x=3/5
 
Kushagra Madhukar
3 years ago
Hello student
The balanced reaction is,
3KMnO4 + 5FeC2O4+ 24H+→ 5Fe3++ 10CO2 + 3Mn2++3K++ 12H2O
Since 3 moles of KMnO4 are required to react with 5 moles of FeC2O4
Hence number of moles of KMnO4 required to react with one mole of FeC2O4 = 3/5 moles
Hope it helps
Regards,
Kushagra